• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Chiranjib

    Log entry time 08:01:16 on August 09, 2006

    Entry number 178594

    keyword=OWL SHIFT SUMMARY

    SHIFT SUMMARY :::

    KIN :: 74.31 P=0.5433



    12:15 A :: ASked for 20 microamps.
    C optics position.

    12:30 A :: 4 cm Dummy.

    the right arm high voltage needed to be reset. ( ALEX did it)

    12:55 A :: No beam. ( MCC needed to perform some tests !)

    1:20 A :: Beam is back. 20 muA.

    1:20 A :: 15 cm Dummy.

    1: 37 A : on Tantalum

    ASked for 12 muA

    1: 53 A : on LH2

    Asked for 20 muA

    2: 03 A : on LH2 ( 20 muA)

    2: 29 A : on LD2 ( 20 muA)

    2: 43 A : on LD2 ( 12 muA)


    The VESDA HALL A alarm went on ( beam way). Called MCC.

    2:50 A : Production run is started

    ~6 A alarm is reset .




    ############################################################################



    RUN SUMMARY :::



    Radiator out.

    12:20 A : RUN 2376 on C Optics with 20 microA.

    12:30 A : RUN 2377 on 4 cm Dummy with 20 microA.

    12:50 A : RUN 2379 on 15 cm Dummy with 20 microA.
    Not completed !! MCC turned the beam off for
    10/15 min.

    1:20 A : Beam is back . 20 muA

    1: 20 A : RUN 2380 on 15 cm Dummy

    1: 37 A : RUN 2381 on Tantalum

    Radiator in ( 5%)

    1: 53 A : RUN 2382 on LH2 ( 12 muA)

    Radiator out

    2: 03 A : RUN 2383 on LH2 ( 20 muA)

    2: 29 A : RUN 2384 on LD2 ( 20 muA)

    Radiator in ( 5%)

    2: 43 A : RUN 2386 on LH2 ( 12 muA)

    2:50 A : Production Run is started

    RUN # Event#

    2388 1.5M
    2389 1.9M
    2390 1.9M
    2391 1.9M
    2392 1.9M
    2393 1.9M
    2394 1.9M
    2395 1.9M
    2396 1.9M
    2397 1.9M
    2398 1.9M
    2399 1.9M
    2400 1.9M
    2401 1.9M
    2402 1.9M
    2403 1.9M
    2404 1.9M
    2405 1.9M
    2406 running....